OpenStudy (anonymous):

When doing integration by parts, what's a good tip to choose which one to be f'(x) and which one g(x)?

OpenStudy (anonymous):

\[\int\limits_{0}^{1}x ^{2}e ^{x}dx\] For example, this one I originally chose f'(x) = x^2 and g(x) = e^x only to find out it doesn't work well

OpenStudy (anonymous):

yeah try integration by parts here \[\int\limits_{0}^{1}x ^{3}e ^{x}\] let u=x^2 and v=e^x

OpenStudy (anonymous):

mark I figured out how to do it thanks though

OpenStudy (anonymous):

thanks mark

OpenStudy (anonymous):

im sorry x=3 and v=e^x

OpenStudy (anonymous):

ok yw, have fun thnx